OpenStudy (gabby1214):

solve a-4=-24 solution is a=?(Integer or simplified fraction)

OpenStudy (benlindquist):

24+4 = a what

OpenStudy (michele_laino):

hint: If I add \(4\) to both sides, I get: \(a-4+4=-24+4\) please simplify